40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6420     Entry time: Wed Mar 14 23:02:09 2012
Author: Koji 
Type: Update 
Category: LSC 
Subject: Locking activity 

Kiwamu and Koji

The target is to realize DRMI or PRMI + one arm with ALS.

The focus of the night is to achive stable lock of the PRMI (SB resonant) with 3f signals.
Particularly, REFL165 is back now, we are aiming to see if any of the 165 signals is useful.

We made a comparison between  REFL33Q/REFL165Q/AS55Q to find any good source of MICH.
However, none of them showed a reasonable shape of the spectra. They don't have reasonable coherence between them.

Nonetheless, we have tried to lock the IFO with those REFL signals. But any of them were useful to keep the PRMI (SB resonant).
The only kind of stable signal for MICH was AS55Q as we could keep the PRMI locked.

ELOG V3.1.3-